# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 3.2 LGW transitions with a gauge-invariant order parameter
The first class of transitions, the LGW transitions, are characterized by the fact that only gauge-invariant matter correlations become critical. Gauge fields have apparently only the role of hindering some scalar degrees of freedom, those that are not gauge invariant, from becoming critical. It seems thus natural to assume that the transition is driven by a gauge-invariant order parameter and that only gauge-invariant matter modes are relevant at criticality. Although this assumption seems natural, it is important to stress that it is not strictly necessary. There are indeed transitions (the LGW× transitions that we will discuss in Sec. 3.3) that are analogous to the LGW transitions as only matter fields are critical, but that require a scalar non-gauge invariant order parameter. Here we consider the conventional case in which the scalar order parameter is gauge invariant. In this case the approach is the conventional one. Once an order parameter has been identified, one considers the corresponding order-parameter field and LGW Lagrangian. The critical behavior is then determined by studying the field-theoretical model as outlined in the previous section 2.2. Note that the gauge fields do not appear in the effective description: the gauge symmetry plays only a role in the choice of the order-parameter field, which should be gauge invariant.
In the next subsections we present some examples of LGW transitions, which will be also useful for the rest of the review. Other examples will be found later in the review.
@@ 3.2.1 Three-dimensional CPN-1 models
The transition in the 3D lattice CPN-1 model, and also the transitions of LAH models (they are discussed in Secs. 5 and 6) in the strong gauge-coupling (small-$\kappa$) regime are examples of LGW transitions. In the CPN-1 field theory the fundamental field is a complex $N$-component unit-length vector ${\bm{z}}({\bm{x}})$, associated with an element of the complex projective manifold CPN-1 [2]. The Lagrangian reads
@@
The stable FP of the RG flow of the LGW theory (10) is expected to control the critical behavior of the continuous transitions in lattice CPN-1 models. In the absence of a stable FP, a first-order transition is expected. [^7] The same effective description applies to antiferromagnetic CPN-1 models, i.e. the model (7) with $J<0$ [note that in the model (9) the sign of $J$ is irrelevant, always corresponding to a ferromagnetic CPN-1 model]. In this case, however, there is an additional ${\mathbb{Z}}_{2}$ symmetry related to translations of one site [207], which implies that the corresponding effective LGW theory (the order-parameter field is related to a staggered gauge-invariant composite operator) must be also invariant under $\Phi\to-\Phi$. This implies the absence of the cubic term in the LGW theory (10), i.e., $w=0$. For $N=3$ this LGW Hamiltonian is equivalent to that of the O(8) vector model, thus predicting a critical behavior belonging to the O(8) vector universality class, with an effective enlargement of the symmetry group of the critical modes from SU(3) to O(8). This prediction has been confirmed numerically [207]. The critical behavior for larger values of $N$ is discussed in Refs. [207, 165].
@@
Numerical simulations of the lattice gauge CPN-1 model (8) confirm the LGW predictions [82]. For $N\geq 3$ the transition is of first order, the first-order nature becoming stronger and stronger as $N$ increases [208]. [^10] It should be noted that results for other models that are expected to behave as the CPN-1 model are less clear. In particular, there is no general agreement that all models undergo a first-order transition for $N=3$, see, e.g., Refs. [64, 209, 210, 211, 207, 165, 82]. In the effective description the gauge group plays no role and thus one expects the same critical behavior in any SU($N$) invariant model that has a bilinear Hermitian order parameter, irrespective of the nature of the gauge group. This conclusion is supported by the results of Ref. [212], which studied a discretized gauge model with Hamiltonian still given by Eq. (8) but with a reduced ${\mathbb{Z}}_{Q}$ gauge symmetry, obtained by requiring the gauge fields to take only the values $\exp(i2\pi n/Q)$ with $n=0,\ldots Q-1$. For $N=2$ the model undergoes an O(3) transition for any $Q\geq 3$, as the CP1 model—this can be shown by explicitly performing the summation over the gauge degrees of freedom—indicating that the nature of the gauge group is not important.
@@
Finally, it is interesting to note that an analysis within the $1/N$-expansion framework indicates that the CPN-1 field theory with Lagrangian (6) is equivalent [64, 2] to the Abelian Higgs field theory (AHFT), or scalar electrodynamics, which will be presented in Sec. 3.4, whose RG flow has a stable charged FP for large values of $N$, see Sec. 4. This result however does not apply to the lattice CPN-1 models. Indeed, gauge fields can be integrated out in the lattice gauge CPN-1 models, obtaining a Hamiltonian that can be written as a sum of local gauge-invariant scalar terms. Thus, gauge modes cannot be relevant for the critical behavior, as instead would be the case if the CPN-1 transition were effectively described by a gauge field theory with dynamical gauge fields, as the AHFT presented in Sec. 3.4 below. A more detailed discussion is presented in Secs. 5 and 6.
@@ 3.2.2 Three-dimensional RPN-1 models
Both models are invariant under SO($N$) global transformations and ${\mathbb{Z}}_{2}$ gauge transformations
${\bm{s}}_{\bm{x}}\to w_{\bm{x}}{\bm{s}}_{\bm{x}},\qquad\sigma_{{\bm{x}},\mu}\to w_{\bm{x}}\sigma_{{\bm{x}},\mu}w_{{\bm{x}}+\hat{\mu}}.$ (13)
@@
Therefore, the LGW field is a symmetric and traceless real tensor $\Phi^{ab}$, obtained by coarse-graining $R^{ab}_{\bm{x}}$, and the LGW Lagrangian [165, 166] is [^12] The LGW Lagrangian (15) is also relevant for the critical behavior of antiferromagnetic RPN-1 models, such as the lattice models (11) with $J<0$, at least for $N\leq 3$. In this case, however, the effective theory is also invariant under $\Phi\to-\Phi$, which implies $w=0$. For $N=3$ the corresponding LGW Hamiltonian is equivalent to that of the O(5) vector theory, implying an effective enlargement of the symmetry group, from O(3) to O(5), at the transition [222]. Numerical results confirm this prediction. For $N\geq 4$ there are discrepancies between the numerical results and the field-theoretical predictions from high-order perturbative analyses of the RG flow of the LGW theory. The origin of these differences is still an open problem, see Ref. [166].
@@
For $N=2$, the field $\Phi^{ab}$ has only two independent components, the $\Phi^{3}$ term is absent, and the Lagrangian (15) is equivalent to that of the standard XY vector model. Thus, continuous transitions should belong to the XY universality class. In this case the field $\Phi^{ab}$ is equivalent to an O(2) vector field. This implies that the RG dimension of $R_{\bm{x}}^{ab}$ coincides with the RG dimension of the vector field in the standard XY model.
@@ 3.2.3 Finite-temperature transition of hadronic matter in the chiral limit
As another example where the LGW scenario is expected to hold, we consider the finite-temperature transition of hadronic matter in the chiral limit, which was analyzed by Pisarski and Wilczek [25] using the LGW approach. The strong hadronic interactions are described by quantum chromodynamics (QCD), which is a non-Abelian gauge theory with gauge group SU(3) and quark fields in the fundamental representation of the gauge group. The thermodynamics of the system is characterized by a finite-temperature transition at $T_{c}\simeq 200$ MeV, which separates a low-temperature confined hadronic phase, in which chiral symmetry is broken, from a high-temperature phase with deconfined quarks and gluons (quark-gluon plasma), in which chiral symmetry is restored [223, 224, 225]. Our present understanding of this finite-$T$ transition is based on the symmetry properties of the system. In the presence of $N_{f}$ light quarks, and, in particular, in their massless limit, the relevant symmetry is the chiral symmetry ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}$, where $L,R$ stand for left and right quark components. Below $T_{c}$, this symmetry is spontaneously broken to SU($N_{f}$)V, the group of vector quark transformations, with a nonzero quark condensate $\langle{\bar{\psi}}\psi\rangle$. The finite-$T$ transition is therefore characterized by the symmetry-breaking pattern ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}\rightarrow{\rm SU}(N_{f})_{V}$. [^13] If the axial U(1)A symmetry is effectively restored at $T_{c}$, the symmetry-breaking pattern becomes ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}\rightarrow{\rm U}(N_{f})_{V}$. The restoration of this symmetry is not expected in QCD. However, semiclassical calculations in the high-temperature phase [24] show that instanton effects are exponentially small for $T\gg T_{c}$, implying that the effect of the breaking the axial U(1)A symmetry becomes small in this limit. This scenario has been also confirmed by numerical studies, see e.g. Refs. [226, 227, 228]. RG studies of the LGW theory appropriate to describe the finite-$T$ transition with symmetry-breaking pattern ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}\rightarrow{\rm U}(N_{f})_{V}$ are reported in Refs. [169, 229].
@@
    where ${\cal L}_{{\rm U}(N_{f})}$ is invariant under ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}$ transformations, and ${\cal L}_{\rm det}$ contains terms involving the determinant of $\Psi_{ij}$, which effectively takes into account the explicit breaking of the U(1)A axial symmetry due to the anomaly [Eq. (16) reports all terms that are relevant for $N_{f}=2$]. Nonvanishing quark masses can be accounted for by adding an external-field term ${\rm Tr}\left(H\Psi+{\rm h.c.}\right)$.
The analysis of the RG flow of the model depends on the value of $N_{f}$. For $N_{f}\geq 4$, a continuous transition is only possible if there is a stable FP of the RG flow that lies in the parameter region in which the symmetry-breaking pattern ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}\rightarrow{\rm SU}(N_{f})_{V}$ occurs. No such FP is found and therefore a first-order transition is predicted. For $N_{f}=3$, the Lagrangian contains a cubic term (the term proportional to $w_{1}$), and therefore the transition is expected to be of first order, too. For $N_{f}=1$ the Lagrangian (16) with $w_{i}=0$ reduces to the O(2) symmetric $\Phi^{4}$ theory, corresponding to the 3D XY universality class, and the term proportional to $w_{1}$ plays the role of an external field. Thus, no transitions are expected, but only a crossover.
@@
We stress that these predictions are based on a number of assumptions, the main one being that the colored gauge modes are effectively decoupled from the critical modes at the transition. If this assumption is correct, then the effective LGW theory in terms of a gauge-invariant bilinear quark operator can provide reliable predictions. See Ref. [225] for a report on the state of art of the lattice numerical results for the chiral transition.
3.3 LGW× transitions with a gauge-dependent order parameter
@@
The simplest examples of these unconventional continuous transitions are those belonging to the 3D O($N$)× (Ising× and XY× for $N=1$ and 2, respectively) universality classes, corresponding to the LGW× transitions associated with the O($N$)-symmetric vector $\Phi^{4}$ theories, cf. Eq. (2). These transitions can be observed in the lattice $N$-vector model minimally coupled with ${\mathbb{Z}}_{2}$-gauge variables [231, 232]. On a cubic lattice its partition function is given by
@@
For $N=1$ there is no global symmetry group, since the global ${\mathbb{Z}}_{2}$ symmetry of the ungauged model has been turned into a local one. However, the global ${\mathbb{Z}}_{2}$ symmetry reemerges in the effective description of the large-$K$ transitions. Their critical features are the same as in the LGW theory associated with the conventional Ising transitions. Therefore, we will name these transitions Ising× transitions.
@@ 3.4 GFT transitions described by effective field theories with gauge fields
LGW and LGW× transitions are both characterized by the fact that the critical behavior is completely determined by the behavior of the scalar variables. Here we consider transitions in which both scalar and gauge correlations are critical at the transition. In this case an appropriate effective field-theory description of the critical behavior requires a gauge field theory (GFT) with the explicit inclusion of gauge fields. Transitions of this type are observed in LAH models, as we shall discuss in Secs. 5 and  6. The corresponding field theory description is provided by the Abelian Higgs field theory (AHFT), or scalar electrodynamics. In the AHFT an $N$-component complex scalar field ${\bm{\Phi}}({\bm{x}})$ is minimally coupled with an electromagnetic field $A_{\mu}({\bm{x}})$. The AHFT Lagrangian reads [17]
@@
It is interesting to note that in lattice gauge models the global symmetry group and the symmetry breaking pattern do not completely characterize the nature of the phase transition. The properties of the gauge modes are crucial to determine the critical behavior. Systems with global SU($N$) symmetry, symmetry-breaking pattern SU($N$$\to$ SU($N-1$) and local U(1) gauge symmetry provide an example of such behavior. If the gauge modes are not critical the transition is of LGW type. The corresponding $\Phi^{4}$ theory is given in Eq. (10), which predicts first-order transitions for $N\geq 3$ and continuous transitions belonging to the O(3) vector universality class for $N=2$. On the other hand, if gauge modes are relevant, the transitions are described by the AHFT (18): they can be continuous for $N>N^{*}\approx 7$, and of first order in the opposite case, and in particular for $N=2$. Therefore, in the two cases the expected behavior at the transition is quite different.

### v3

@@ 3.2 LGW transitions with a gauge-invariant order parameter
The first class of transitions, the LGW transitions, is characterized by the fact that only gauge-invariant matter correlations are critical. Gauge fields have apparently only the role of hindering some scalar degrees of freedom, those that are not gauge invariant, from becoming critical. It seems thus natural to assume that the transition is driven by a gauge-invariant order parameter and that only gauge-invariant matter modes are relevant at criticality. Although this assumption seems natural, it is important to stress that it is not strictly necessary. There are indeed transitions (the LGW× transitions that we will discuss in Sec. 3.3) that are analogous to the LGW transitions as only matter fields are critical, but that require a scalar nongauge invariant order parameter. If the scalar order parameter is gauge invariant, the approach is the conventional one. Once an order parameter has been identified, one considers the corresponding order-parameter field and LGW Lagrangian. The critical behavior is then determined by studying the field-theoretical model as outlined in Sec. 2.2. Note that the gauge fields do not appear in the effective description: the gauge symmetry plays only a role in selecting the order-parameter field, which should be gauge invariant.
In the next subsections we present some examples of LGW transitions, which will be also useful for the remainder of the review. Other examples will be presented later in the review.
@@ 3.2.1 Three-dimensional CPN-1 models
The transition in the 3D lattice CPN-1 model, and also the transitions in LAH models (they are discussed in Secs. 5 and 6) in the strong gauge-coupling (small-$\kappa$) regime, are examples of LGW transitions. In the CPN-1 field theory the fundamental field is a complex $N$-component unit-length vector ${\bm{z}}({\bm{x}})$, associated with an element of the complex projective manifold CPN-1 [2]. The Lagrangian reads
@@
The stable FP of the RG flow of the LGW theory (10) is expected to control the critical behavior of the continuous transitions in lattice CPN-1 models. In the absence of a stable FP, a first-order transition is expected. [^7] The same effective description applies to antiferromagnetic CPN-1 models, i.e., to the model with Hamiltonian (7) and $J<0$ (in the Hamiltonian (8) the sign of $J$ is irrelevant, and therefore the model is always ferromagnetic). In this case, however, there is an additional ${\mathbb{Z}}_{2}$ symmetry related to translations of one site [236], which implies that the corresponding effective LGW theory (the order-parameter field is related to a staggered gauge-invariant composite operator) must also be invariant under $\Phi\to-\Phi$. This implies the absence of the cubic term in the LGW theory (10), i.e., $w=0$. For $N=3$ this LGW Hamiltonian is equivalent to that of the O(8) vector model, thus predicting a critical behavior belonging to the O(8) vector universality class, with an effective enlargement of the symmetry group of the critical modes from SU(3) to O(8). This prediction has been confirmed numerically [236]. The critical behavior for larger values of $N$ is discussed in Refs. [236, 196].
@@
Numerical simulations of the lattice gauge CPN-1 model (8) confirm the LGW predictions [82]. For $N\geq 3$ the transition is of first order, the first-order nature becoming stronger and stronger as $N$ increases [237]. [^10] It should be noted that results for other models that are expected to behave as the CPN-1 model are less clear. In particular, there is no general agreement that all models undergo a first-order transition for $N=3$, see, e.g., Refs. [64, 238, 239, 240, 236, 196, 82]. In the effective description the gauge group plays no role and thus one expects the same critical behavior in any SU($N$) invariant model that has a bilinear Hermitian order parameter, irrespective of the nature of the gauge group. This conclusion is supported by the results of Ref. [241], which studied a discretized gauge model with Hamiltonian still given by Eq. (8) but with a reduced ${\mathbb{Z}}_{Q}$ gauge symmetry, obtained by requiring the gauge fields to take only the values $\exp(2\pi in/Q)$ with $n=0,\ldots Q-1$. For $N=2$ the model undergoes an O(3) transition for any $Q\geq 3$, as the CP1 model—this can be shown by explicitly performing the summation over the gauge degrees of freedom—indicating that the nature of the gauge group is not important.
@@
Finally, it is interesting to note that an analysis in the $1/N$-expansion framework indicates that the CPN-1 field theory with Lagrangian (6) is equivalent [64, 2] to the Abelian Higgs field theory (AHFT), or scalar electrodynamics, which will be presented in Sec. 3.4, whose RG flow has a stable charged FP for large values of $N$, see Sec. 4. This result however does not apply to the lattice CPN-1 models. Indeed, gauge fields can be integrated out in the lattice gauge CPN-1 models, obtaining a Hamiltonian that can be written as a sum of local gauge-invariant scalar terms. Thus, gauge modes cannot be relevant for the critical behavior, as instead would be the case if the CPN-1 transition were effectively described by a gauge field theory with dynamical gauge fields, as the AHFT presented in Sec. 3.4 below. A more detailed discussion is presented in Secs. 5 and 6.
@@ 3.2.2 Three-dimensional RPN-1 models
Both models are invariant under SO($N$) global transformations and the ${\mathbb{Z}}_{2}$ gauge transformations
${\bm{s}}_{\bm{x}}\to w_{\bm{x}}{\bm{s}}_{\bm{x}},\qquad\sigma_{{\bm{x}},\mu}\to w_{\bm{x}}\sigma_{{\bm{x}},\mu}w_{{\bm{x}}+\hat{\mu}},\qquad w_{{\bm{x}}}=\pm 1.$ (13)
@@
Therefore, the LGW field is a symmetric and traceless real tensor $\Phi^{ab}({\bm{x}})$, obtained by coarse-graining $R^{ab}_{\bm{x}}$, and the LGW Lagrangian [196, 197] is [^12] The LGW Lagrangian (15) is also relevant for the critical behavior of antiferromagnetic RPN-1 models, such as the lattice models (11) with $J<0$, at least for $N\leq 3$. In this case, however, the effective theory is also invariant under $\Phi\to-\Phi$, which implies $w=0$. For $N=3$ the corresponding LGW Hamiltonian is equivalent to that of the O(5) vector theory, implying an effective enlargement of the symmetry group, from O(3) to O(5), at the transition [251]. Numerical results confirm this prediction. For $N\geq 4$ there are discrepancies between the numerical results and the field-theoretical predictions obtained from high-order perturbative analyses of the RG flow of the LGW theory. The origin of these differences is still an open problem, see Refs. [197] and [252].
@@
For $N=2$ the field $\Phi^{ab}$ has only two independent components, the $\Phi^{3}$ term is absent, and the Lagrangian (15) is equivalent to that of the standard XY vector model. Thus, continuous transitions should belong to the XY universality class. In this case the field $\Phi^{ab}$ is equivalent to an O(2) vector field. This implies that the RG dimension of $R_{\bm{x}}^{ab}$ coincides with the RG dimension of the vector field in the standard XY model.
@@ 3.2.3 Finite-temperature transition of hadronic matter in the chiral limit
As another example where the LGW approach is expected to hold, we consider the finite-temperature transition of hadronic matter in the chiral limit, which was analyzed by Pisarski and Wilczek [25] using the LGW approach. The strong hadronic interactions are described by quantum chromodynamics (QCD), which is a non-Abelian gauge theory with gauge group SU(3) and quark fields in the fundamental representation of the gauge group. The thermodynamics of the system is characterized by a finite-temperature transition at $T_{c}\simeq 200$ MeV, which separates a low-temperature confined hadronic phase, in which chiral symmetry is broken, from a high-temperature phase with deconfined quarks and gluons (quark-gluon plasma), in which chiral symmetry is restored [253, 254, 255]. Our present understanding of this finite-$T$ transition is based on the symmetry properties of the system. In the presence of $N_{f}$ light quarks, and, in particular, in their massless limit, the relevant symmetry is the chiral symmetry ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}$, where $L,R$ stand for left and right quark components. Below $T_{c}$, this symmetry is spontaneously broken to SU($N_{f}$)V, the group of vector quark transformations, with a nonzero quark condensate $\langle{\bar{\psi}}\psi\rangle$. The finite-$T$ transition is therefore characterized by the symmetry-breaking pattern ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}\rightarrow{\rm SU}(N_{f})_{V}$. [^13] If the axial U(1)A symmetry is effectively restored at $T_{c}$, the symmetry-breaking pattern becomes ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}\rightarrow{\rm U}(N_{f})_{V}$. The restoration of this symmetry is not expected in QCD. However, semiclassical calculations in the high-temperature phase [24] show that instanton effects are exponentially small for $T\gg T_{c}$, implying that the effect of the breaking the axial U(1)A symmetry becomes small in this limit. This scenario has been also confirmed by numerical studies, see, e.g., Refs. [256, 257, 258]. RG studies of the LGW theory appropriate to describe the finite-$T$ transition with symmetry-breaking pattern ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}\rightarrow{\rm U}(N_{f})_{V}$ are reported in Refs. [200, 259].
@@
    where ${\cal L}_{{\rm U}(N_{f})}$ is invariant under ${\rm U}(N_{f})_{L}\otimes{\rm U}(N_{f})_{R}$ transformations. The Lagrangian ${\cal L}_{\rm det}$ contains terms involving the determinant of $\Psi_{ij}$, which effectively takes into account the explicit breaking of the U(1)A axial symmetry due to the anomaly (Eq. (16) reports all terms that are relevant for $N_{f}=2$). Nonvanishing quark masses can be accounted for by adding an external-field term ${\rm Tr}\left(H\Psi+{\rm h.c.}\right)$.
The analysis of the RG flow of the model depends on the value of $N_{f}$. For $N_{f}\geq 4$, a continuous finite-$T$ transition is only possible if there is a stable FP of the RG flow that lies in the parameter region in which the symmetry-breaking pattern ${\rm SU}(N_{f})_{L}\otimes{\rm SU}(N_{f})_{R}\rightarrow{\rm SU}(N_{f})_{V}$ occurs. Since there is no FP associated with this symmetry-breaking pattern, continuous transitions cannot occur (thus, contradicting the first assumption leading to the LGW approach), and therefore first-order transitions are generally predicted. For $N_{f}=3$, the Lagrangian contains a cubic term (the term proportional to $w_{1}$), and therefore the transition is expected to be of first order, too. For $N_{f}=1$ the Lagrangian (16) with $w_{i}=0$ reduces to the O(2) symmetric $\Phi^{4}$ theory, corresponding to the 3D XY universality class, and the term proportional to $w_{1}$ plays the role of an external field. Thus, no transitions are expected, but only a crossover.
@@
We stress that these predictions are based on a number of assumptions, the main one being that the colored gauge modes are effectively decoupled from the critical modes at the transition. If this assumption is correct, then the effective LGW theory in terms of a gauge-invariant bilinear quark operator can provide reliable predictions. See, e.g., Refs. [255, 261] for a report on the lattice numerical results obtained for the chiral transition.
3.3 ${}^{\,\times}$ transitions with a gauge-dependent order parameter
@@
3D O($N$)× (Ising× and XY× for $N=1$ and 2, respectively) transitions are the simplest examples of this type of unconventional transitions. They admit an effective description in terms of the O($N$)-symmetric $\Phi^{4}$ theory with Lagrangian  (2) and can be observed, e.g., in the lattice $N$-vector model minimally coupled with ${\mathbb{Z}}_{2}$-gauge variables [262, 263]. On a cubic lattice its partition function is given by
@@
For $N=1$ there is no global symmetry group, since the global ${\mathbb{Z}}_{2}$ symmetry of the ungauged model has been turned into a local one. However, the global ${\mathbb{Z}}_{2}$ symmetry reemerges in the effective description of the large-$K$ transitions. Indeed, their critical features are the same as in the conventional Ising transition. Therefore, we will name these transitions Ising× transitions.
@@ 3.4 GFT transitions described by effective field theories with gauge fields
LGW and LGW× transitions are both characterized by the fact that the critical behavior is completely determined by the behavior of the scalar variables. Here we consider transitions in which both scalar and gauge correlations are critical at the transition. In this case an appropriate effective field-theory description of the critical behavior requires a gauge field theory (GFT) with the explicit inclusion of gauge fields. Transitions of this type are observed in LAH models, as we shall discuss in Secs. 5 and 6. The corresponding field theory description is provided by the Abelian Higgs field theory (AHFT), or scalar QED. In the AHFT an $N$-component complex scalar field ${\bm{\Phi}}({\bm{x}})$ is minimally coupled with an electromagnetic field $A_{\mu}({\bm{x}})$. The AHFT Lagrangian reads [17]
@@
It is interesting to note that in lattice gauge models the global symmetry group and the symmetry breaking pattern do not completely characterize the nature of the phase transition. The properties of the gauge modes are crucial to determine the critical behavior. Systems with global SU($N$) symmetry, symmetry-breaking pattern SU($N$$\to$ SU($N-1$), and local U(1) gauge symmetry provide an example of such behavior. If the gauge modes are not critical the transition is of LGW type. The corresponding $\Phi^{4}$ theory is given in Eq. (10), which predicts first-order transitions for $N\geq 3$ and continuous transitions belonging to the O(3) vector universality class for $N=2$. On the other hand, if gauge modes are relevant, the transitions are described by the AHFT (18): they can be continuous for $N>N^{*}\approx 7$, and of first order in the opposite case, and in particular for $N=2$. Therefore, in the two cases the expected behavior at the transition is quite different.